OpenStudy (anonymous):

If 3 dice are rolled at the same time, what is the probability of all 3 dice showing even face values showing 2, 4, or 6?

OpenStudy (kropot72):

The probability of each of the dice showing an even face value is 1/2. The result of rolling each die has no effect on the result of rolling the others, therefore the outcomes are independent. The required probability is therefore \[P(all\ 3\ showing\ even\ face\ value)=\frac{1}{2}\times\frac{1}{2}\times\frac{1}{2}\]